OpenStudy (jenniferjuice):

The sum of two consecutive even integers is 226. What is the larger integer?

OpenStudy (whpalmer4):

One even integer can be written as 2n, where n is an integer. The next consecutive even integer after 2n would be 2n + 2, right? So 2n + (2n+2) = 226. Find n, then use it to find the value of 2n+2.

OpenStudy (jenniferjuice):

ummm so do i ....

OpenStudy (whpalmer4):

you solve 2n + (2n+2) = 226...

OpenStudy (whpalmer4):

\[2n+(2n+2)=226\]\[4n+2=226\]your turn...

OpenStudy (jenniferjuice):

alright let me write this down

OpenStudy (jenniferjuice):

and try to solve it

OpenStudy (anonymous):

i have an idea, lets guess and check

OpenStudy (whpalmer4):

operators are standing by to check your answer :-)

OpenStudy (anonymous):

\(110+112=222\) nope

OpenStudy (anonymous):

bet i can guess it on the second try though...

OpenStudy (whpalmer4):

or even on the first, if you tried the two numbers surrounding half of 226 :-)

OpenStudy (jenniferjuice):

would i add 2n +2n? first?

OpenStudy (anonymous):

i am not that smart but if you order now you will get two plus the can opener

OpenStudy (anonymous):

@jenniferjuice you are smarter than that no need to solve a silly equation just guess and check

OpenStudy (whpalmer4):

\[4n + 2 = 226\]subtract 2 from both sides\[4n=224\]\[n=56\]so the big number was 2n+2 = 2(56)+2 =

OpenStudy (jenniferjuice):

oh i thought about that as well but i wasn't quite sure

OpenStudy (anonymous):

like i said, since \(110+112\) is too small, try a slightly larger one

OpenStudy (jenniferjuice):

@whpalmer4